There are too conductors A & B of the same material.

having length l & 2l: & having radii r & r/2 respectively. what is the ratio of their resistances?


1
Expert's answer
2021-06-01T14:53:45-0400

Gives

"l_1=L"

"l_2=2L"

"r_1=r"

"r_2=\frac{r}{2}"

We know that resistance


"R=\frac{\rho l}{A}"

Now

"\frac{R_1}{R_2}=\frac{l_1}{l_2}\times (\frac{r_2}{r_1})^2"

Put Value

"\frac{R_1}{R_2}=\frac{L}{2L}\times (\frac{\frac{r}{2}}{r})^2"

"\frac{R_1}{R_2}=\frac{1}{8}"